# Jiaming Wei    Dihan Zou

### v2

@@ 3.2 Intervention as the optimal regulation
Suppose condition (LF) is violated, $f$ is nonincreasing and log-concave, and the demand function $P^{-1}$ is strictly log-concave. [^18] These conditions complement those in AmadorBagwell22Regulation (Proposition 1, p. 1735) and hold for a broad class of demand and cost specifications. For example, if demand is microfounded by a continuum of consumers with i.i.d. valuations drawn from $G$, strict log-concavity of the density $g$ implies $P^{-1}(p)=1-G(p)$ is strictly log-concave. Also, log-concavity of $P^{-1}$ is stronger than the strict concavity of the revenue function $qP(q)$, so our Assumption 1 remains valid (see ZouLogConcave). Then there exists $\overline{k}>0$ such that for all $k\in[0,\overline{k})$, the optimal regulation is a progressive price cap $\tau_{\hat{p}}(p)$. Moreover, there exists a cutoff $\hat{c}\in(0,\overline{c}_{LF})$ such that firms with marginal cost above (resp. below) $\hat{c}$ choose prices above (resp. below) $\hat{p}$.
The proposition characterizes the optimal regulatory instrument, the induced pricing behavior, and the resulting allocation. When intervention is warranted, the optimal policy consists of both delegation and taxation. Higher prices are discouraged through increasing tax rates, and firms self-select into the two regions according to cost. High-cost types ($c>\hat{c}$) set $p>\hat{p}$ and are taxed; sufficiently high-cost types may face prohibitive taxes and be excluded. Low-cost types ($c\leq\hat{c}$) set $p=\hat{p}$ and are untaxed; [^19] It is well-known in the mechanism literature that constraints on transfers can lead to pooling allocation. See, e.g., KrishnaMorgan08, PaiVohra14. those with sufficiently low costs may instead optimally charge their laissez-faire price $P(q_{LF}(c))<\hat{p}$. Lemma B.8 reports explicit expressions for firms’ optimal pricing and the resulting tax schedule.
@@
Now consider $\alpha>0$, so the regulator also cares about (c) firm profit. Distortionary taxation now reduces not only consumer surplus but also the firm’s gains from trade. Regulation must therefore navigate a three-way tension among access, affordability, and profitability. [^20] As emphasized by GuoShmaya25RobustRegu, the regulator must balance: mitigating underproduction (access and profitability), protecting consumer surplus (affordability), and limiting overproduction (irrelevant under no-subsidy constraints).
@@
Figure 3: Firm’s inverse demand function $P_{\tau}(q)$ (left) and the induced pricing strategy $p^{*}(c)$ (right) with flexible transfers according to BaronMyerson82. P(q)=1qP(q)=1-q, $c\sim U[0,1]$, $\alpha=1$.
@@ 4 Conclusion
Constraints on policy instruments have been recognized for decades as posing important practical and theoretical challenges, yet they have received relatively limited formal treatment. In their classic textbook (p. 155), LaffontTiroleBook93 note one such challenge:
@@ Appendix B Proof of Proposition 1 and 2
then the problem is equivalent to that of AmadorBagwell22Regulation. Methodologically, we use Pontryagin approach to a mechanism problem with type-dependent constraints (e.g., XiaoPontryagin; Xiao25JMP).
@@
$\max_{\Pi,q,z}\int_{0}^{1}\left[V(q(c))-cq(c)-(1-\alpha)\Pi(c)\right]f(c)dc$ (6)
@@
We show that nondecreasing $q(\cdot)$ will lead to a contradiction following LaffontTiroleBook93 (p.157). Condition (11) is equivalent to
@@
Note that the first-order condition is equivalent to the envelope condition, implying the latter is also sufficient and necessary for optimality. We will use the envelope condition in our analysis for convenience. Recall that $\Pi(c)=\max_{x}r(x)-cq(x)-k\cdot\mathbbm{1}_{q(x)>0}$. Since the maximand is affine (and thus convex) in $c$, $\Pi(c)$ is a convex function, which implies it is continuous and almost everywhere differentiable. By envelope theorem (e.g., MilgromSegal02Env), $\Pi^{\prime}(c)=-q(c)$ wherever differentiable. Therefore, a set of sufficient and necessary conditions for global (IC) is that (i) $\Pi(c)$ is continuous and nonnegative; (ii) $\Pi^{\prime}(c)=-q(c)$ wherever differentiable; and (iii) $q(c)$ is nonincreasing. We summarize these condition concisely in the form of “law of motion”, (7)(8), in the optimal control problem.
@@
Lemma B.2 (AmadorBagwell22Regulation).
@@
$H=[V(q)-cq-(1-\alpha)\Pi]f-\lambda_{\Pi}q+\lambda_{q}z.$
@@
where $\gamma(c)\geq 0$ is some piecewise continuous function of $c$. By Theorem 5.1 in SeierstadSydsaeter93OptCtrl (pp. 317-319), we derive the following sufficient and necessary conditions [^25] The necessity is guaranteed according to pp. 335 - 336 of SeierstadSydsaeter93OptCtrl. The variable $\beta$ in (12) and (13) is a scale parameter for the jump of costate at $c=0$. By conditions (5.37) - (5.38) of SeierstadSydsaeter93OptCtrl, the costate cannot jump at $c\in(0,1]$. Moreover, the no-subsidy constraint (9) being slack at $c=0$ or $q^{\prime}(0)$ not existing implies $\beta=0$. In condition (12), $\lambda_{\Pi}(0_{+})=-\beta\frac{\partial g(q,\Pi,0)}{\partial\Pi}=\beta\geq 0$. For condition (13) to be consistent with the primal condition (14), we need $\lambda_{q}(0_{+})=-\beta\frac{\partial g(q,\Pi,c)}{\partial q}|_{c=0}\geq 0$, which imposes a requirement on $q$, i.e., $q(0)\geq q_{LF}(0)$. from the Lagrangian: $\forall c\in[0,1]$
@@
Note that we allow for discontinuity of $q(c)$ and the corresponding condition is in (14) (see XiaoPontryagin (p. 7)). We are ready to prove Proposition 1, which provides a simple condition for the scope of intervention.
@@
Next, we turn to the optimal policy when condition (LF) fails. We pin down the behavior of the optimal policy by exploiting the necessity of (12) - (15). Recall that Lemma B.2 shows that $q(\cdot)$ can jump at $\overline{c}$. The following Lemma B.3 shows that $q(\cdot)$ does not have other discontinuity than $\overline{c}$. This is consistent with the intuition of continuous policies in BaronMyerson82. A jump at any non-excluded type is either not incentive compatible, or not optimally extracting the firm’s rent. Strict concavity of the market revenue function (Assumption 1) and continuous density function $f$ of types constitute sufficient conditions to avoid interior jumps for our policy.
@@
Without loss, we can restrict attention to left- and right- continuous $q(c)$, because single-point jumps account to zero measure. We first show that $q(c)$ has no jump for interior $c\in(0,\overline{c})$. Our argument follows MorrisShadmehr23Inspiring, which dates back to Arrow666. Suppose there is a jump of $q(c)$ at some $\sigma\in(0,\overline{c})$. Since $q(c)$ is decreasing, we must have $q(\sigma_{+})<q(\sigma_{-})$, where $q(\sigma_{+}):=\lim_{c\rightarrow\sigma_{+}}q(c)$, and $q(\sigma_{-}):=\lim_{c\rightarrow\sigma_{-}}q(c)$.
@@
On the other hand, $\lambda_{q}(\sigma)=0$ at the jump point of its corresponding state $q(\cdot)$. [^26] See condition (3.74) of Theorem 3.7 in Chapter 3, SeierstadSydsaeter93OptCtrl (p. 196-198). Also see condition (4.13) in Hellwig10ECMA and XiaoPontryagin (p. 7). Since $\lambda_{q}(\sigma)\geq 0$ by the primal condition (14), we have
@@
We first show existence by invoking Theorem 5.5 of SeierstadSydsaeter93OptCtrl, and check the problem (6) - (10) satisfies their conditions. Denote the pointwise welfare as
$W(q,\Pi,c):=V(q)-cq-(1-\alpha)\Pi$
@@
is a rectangle unbounded on one side, which is convex. Therefore, all conditions of Theorem 5.5 of SeierstadSydsaeter93OptCtrl hold. As a result, there exists an optimal solution $(q^{*},\Pi^{*},z^{*})$ for the optimal control problem (6) - (10).
For uniqueness, we follow the approach of KangWatt24.272727We also note that unlike KangWatt24, our regulation problem is a convex program with non-majorization constraints: $q(c)P(q(c))-cq(c)-k\geq\int_{c}^{1}q(x)dx.$ Therefore, we cannot proceed using the novel approach provided in KangWatt24, neither can we adopt the approach of KleinerMajorizationECMA. Suppose there are two distinct optimal mechanisms $(q_{1},\Pi_{1})$ and $(q_{2},\Pi_{2})$ that differ for some positive-measure range of costs. Then each of the mechanism must satisfy IC and the no-subsidy constraint $qP(q)-cq\geq\Pi$. Consider an alternative mechanism $(\frac{q_{1}+q_{2}}{2},\frac{\Pi_{1}+\Pi_{2}}{2})$. The new mechanism satisfies IC because of linearity. It also satisfies the no-subsidy constraint because $qP(q)$ is strictly concave, we have for each $c$
@@
The first two terms $(1-\alpha)\frac{F(c)}{f(c)}+\frac{\Gamma(\hat{c})+\beta}{f(c)}$ is nondecreasing as $f(c)$ is nonincreasing. What remains is to show $\frac{\int_{c}^{1}P^{-1}[\phi(x;\Gamma(\hat{c}))]\,dx+k}{P^{-1}[\phi(c;\Gamma(\hat{c}))]}$ is decreasing. Denote the demand function to be $D:=P^{-1}$, and define the composite function $H:=D\circ\phi_{\Gamma(\hat{c})}$. It is equivalent to show $\frac{\int_{c}^{1}H(x)\,dx+k}{H(c)}$ is decreasing. First, consider the case $k=0$, then $\frac{\int_{c}^{1}H(x)\,dx}{H(c)}$ is strictly decreasing if $H$ is strictly log-concave. [^28] See Theorem 6 in BagBerg05Logconcave. Next, we show that a set of sufficient conditions for $H$ to be strictly log-concave is that $D$ is strictly log-concave, as well as $f$ being log-concave and nonincreasing. By Theorem 3 in ZouLogConcave (p. 5), a set of sufficient conditions for the composite function $H=D\circ\phi_{\Gamma(\hat{c})}$ to be strictly log-concave is (i) $D$ is strictly log-concave and decreasing; (ii) $\phi_{\Gamma(\hat{c})}$ is monotone and convex. We only need to show that $f$ being log-concave and decreasing leads to convex $\phi_{\Gamma(\hat{c})}$, which is true by directly verifying the second-order derivative:

### v3

@@ 3.2 Intervention as the optimal regulation
Suppose condition (LF) is violated, $f$ is nonincreasing and log-concave, and the demand function $P^{-1}$ is strictly log-concave. [^18] These conditions complement those in Amador and Bagwell (2022) (Proposition 1, p. 1735) and hold for a broad class of demand and cost specifications. For example, if demand is microfounded by a continuum of consumers with i.i.d. valuations drawn from $G$, strict log-concavity of the density $g$ implies $P^{-1}(p)=1-G(p)$ is strictly log-concave. Also, log-concavity of $P^{-1}$ is stronger than the strict concavity of the revenue function $qP(q)$, so our Assumption 1 remains valid (see Zou (2025)). Then there exists $\overline{k}>0$ such that for all $k\in[0,\overline{k})$, the optimal regulation is a progressive price cap $\tau_{\hat{p}}(p)$. Moreover, there exists a cutoff $\hat{c}\in(0,\overline{c}_{LF})$ such that firms with marginal cost above (resp. below) $\hat{c}$ choose prices above (resp. below) $\hat{p}$.
The proposition characterizes the optimal regulatory instrument, the induced pricing behavior, and the resulting allocation. When intervention is warranted, the optimal policy consists of both delegation and taxation. Higher prices are discouraged through increasing tax rates, and firms self-select into the two regions according to cost. High-cost types ($c>\hat{c}$) set $p>\hat{p}$ and are taxed; sufficiently high-cost types may face prohibitive taxes and be excluded. Low-cost types ($c\leq\hat{c}$) set $p=\hat{p}$ and are untaxed; [^19] It is well-known in the mechanism design literature that constraints on transfers can lead to pooling allocation. See, e.g., Krishna and Morgan (2008), Pai and Vohra (2014). those with sufficiently low costs may instead optimally charge their laissez-faire price $P(q_{LF}(c))<\hat{p}$. Lemma B.8 reports explicit expressions for firms’ optimal pricing and the resulting tax schedule.
@@
Now consider $\alpha>0$, so the regulator also cares about (c) firm profit. Distortionary taxation now reduces not only consumer surplus but also the firm’s gains from trade. Regulation must therefore navigate a three-way tension among access, affordability, and profitability. [^20] As emphasized by Guo and Shmaya (2025), the regulator must balance: mitigating underproduction (access and profitability), protecting consumer surplus (affordability), and limiting overproduction (irrelevant under no-subsidy constraints).
@@
Figure 3: Firm’s inverse demand function $P_{\tau}(q)$ (left) and the induced pricing strategy $p^{*}(c)$ (right) with flexible transfers according to Baron and Myerson (1982). P(q)=1qP(q)=1-q, $c\sim U[0,1]$, $\alpha=1$.
@@ 4 Conclusion
Constraints on policy instruments have been recognized for decades as posing important practical and theoretical challenges, yet they have received relatively limited formal treatment. In their classic textbook (p. 155), Laffont and Tirole (1993) note one such challenge:
@@ Appendix B Proof of Proposition 1 and 2
then the problem is equivalent to that of Amador and Bagwell (2022). Methodologically, we use Pontryagin approach to a mechanism problem with type-dependent constraints (e.g., Xiao (2025a); Xiao (2025b)).
@@
$\max_{\Pi,q,z}\int_{0}^{1}\left[V(q(c))-cq(c)-k\cdot\mathbbm{1}_{q(c)>0}-(1-\alpha)\Pi(c)\right]f(c)dc$ (6)
@@
We show that nondecreasing $q(\cdot)$ will lead to a contradiction following Laffont and Tirole (1993) (p.157). Condition (11) is equivalent to
@@
Note that the first-order condition is equivalent to the envelope condition, implying the latter is also sufficient and necessary for optimality. We will use the envelope condition in our analysis for convenience. Recall that $\Pi(c)=\max_{x}r(x)-cq(x)-k\cdot\mathbbm{1}_{q(x)>0}$. Since the maximand is affine (and thus convex) in $c$, $\Pi(c)$ is a convex function, which implies it is continuous and almost everywhere differentiable. By envelope theorem (e.g., Milgrom and Segal (2002)), $\Pi^{\prime}(c)=-q(c)$ wherever differentiable. Therefore, a set of sufficient and necessary conditions for global (IC) is that (i) $\Pi(c)$ is continuous and nonnegative; (ii) $\Pi^{\prime}(c)=-q(c)$ wherever differentiable; and (iii) $q(c)$ is nonincreasing. We summarize these condition concisely in the form of “law of motion”, (7)(8), in the optimal control problem.
@@
Lemma B.2 (Amador and Bagwell (2022)).
@@
$H=[V(q)-cq-(1-\alpha)\Pi-k]f-\lambda_{\Pi}q+\lambda_{q}z.$
@@
where $\gamma(c)\geq 0$ is some piecewise continuous function of $c$. By Theorem 5.1 in Seierstad and Sydsæster (1993) (pp. 317-319), we derive the following sufficient and necessary conditions [^25] The necessity is guaranteed according to pp. 335 - 336 of Seierstad and Sydsæster (1993). The variable $\beta$ in (12) and (13) is a scale parameter for the jump of costate at $c=0$. By conditions (5.37) - (5.38) of Seierstad and Sydsæster (1993), the costate cannot jump at $c\in(0,1]$. Moreover, the no-subsidy constraint (9) being slack at $c=0$ or $q^{\prime}(0)$ not existing implies $\beta=0$. In condition (12), $\lambda_{\Pi}(0_{+})=-\beta\frac{\partial g(q,\Pi,0)}{\partial\Pi}=\beta\geq 0$. For condition (13) to be consistent with the primal condition (14), we need $\lambda_{q}(0_{+})=-\beta\frac{\partial g(q,\Pi,c)}{\partial q}|_{c=0}\geq 0$, which imposes a requirement on $q$, i.e., $q(0)\geq q_{LF}(0)$. from the Lagrangian: $\forall c\in[0,1]$
@@
Note that we allow for discontinuity of $q(c)$ and the corresponding condition is in (14) (see Xiao (2025a) (p. 7)). We are ready to prove Proposition 1, which provides a simple condition for the scope of intervention.
@@
Next, we turn to the optimal policy when condition (LF) fails. We pin down the behavior of the optimal policy by exploiting the necessity of (12) - (15). Recall that Lemma B.2 shows that $q(\cdot)$ can jump at $\overline{c}$. The following Lemma B.3 shows that $q(\cdot)$ does not have other discontinuity than $\overline{c}$. This is consistent with the intuition of continuous policies in Baron and Myerson (1982). A jump at any non-excluded type is either not incentive compatible, or not optimally extracting the firm’s rent. Strict concavity of the market revenue function (Assumption 1) and continuous density function $f$ of types constitute sufficient conditions to avoid interior jumps for our policy.
@@
Without loss, we can restrict attention to left- and right- continuous $q(c)$, because single-point jumps account to zero measure. We first show that $q(c)$ has no jump for interior $c\in(0,\overline{c})$. Our argument follows Morris and Shadmehr (2023), which dates back to Arrow (1966). Suppose there is a jump of $q(c)$ at some $\sigma\in(0,\overline{c})$. Since $q(c)$ is decreasing, we must have $q(\sigma_{+})<q(\sigma_{-})$, where $q(\sigma_{+}):=\lim_{c\rightarrow\sigma_{+}}q(c)$, and $q(\sigma_{-}):=\lim_{c\rightarrow\sigma_{-}}q(c)$.
@@
On the other hand, $\lambda_{q}(\sigma)=0$ at the jump point of its corresponding state $q(\cdot)$. [^26] See condition (3.74) of Theorem 3.7 in Chapter 3, Seierstad and Sydsæster (1993) (p. 196-198). Also see condition (4.13) in Hellwig (2010) and Xiao (2025a) (p. 7). Since $\lambda_{q}(\sigma)\geq 0$ by the primal condition (14), we have
@@
We first show existence by invoking Theorem 5.5 of Seierstad and Sydsæster (1993), and check the problem (6) - (10) satisfies their conditions. Denote the pointwise welfare as
$W(q,\Pi,c):=V(q)-cq-k-(1-\alpha)\Pi$
@@
is a rectangle unbounded on one side, which is convex. Therefore, all conditions of Theorem 5.5 of Seierstad and Sydsæster (1993) hold. As a result, there exists an optimal solution $(q^{*},\Pi^{*},z^{*})$ for the optimal control problem (6) - (10).
For uniqueness, we follow the approach of Kang and Watt (2024). [^27] We also note that unlike Kang and Watt (2024), our regulation problem is a convex program with non-majorization constraints: $q(c)P(q(c))-cq(c)-k\geq\int_{c}^{1}q(x)dx.$ Therefore, we cannot proceed using the novel approach provided in Kang and Watt (2024), neither can we adopt the approach of Kleiner et al. (2021). Suppose there are two distinct optimal mechanisms $(q_{1},\Pi_{1})$ and $(q_{2},\Pi_{2})$ that differ for some positive-measure range of costs. Then each of the mechanism must satisfy IC and the no-subsidy constraint $qP(q)-cq\geq\Pi$. Consider an alternative mechanism $(\frac{q_{1}+q_{2}}{2},\frac{\Pi_{1}+\Pi_{2}}{2})$. The new mechanism satisfies IC because of linearity. It also satisfies the no-subsidy constraint because $qP(q)$ is strictly concave, we have for each $c$
@@
The first two terms $(1-\alpha)\frac{F(c)}{f(c)}+\frac{\Gamma(\hat{c})+\beta}{f(c)}$ is nondecreasing as $f(c)$ is nonincreasing. What remains is to show $\frac{\int_{c}^{1}P^{-1}[\phi(x;\Gamma(\hat{c}))]\,dx+k}{P^{-1}[\phi(c;\Gamma(\hat{c}))]}$ is decreasing. Denote the demand function to be $D:=P^{-1}$, and define the composite function $H:=D\circ\phi_{\Gamma(\hat{c})}$. It is equivalent to show $\frac{\int_{c}^{1}H(x)\,dx+k}{H(c)}$ is decreasing. First, consider the case $k=0$, then $\frac{\int_{c}^{1}H(x)\,dx}{H(c)}$ is strictly decreasing if $H$ is strictly log-concave. [^28] See Theorem 6 in Bagnoli and Bergstrom (2005). Next, we show that a set of sufficient conditions for $H$ to be strictly log-concave is that $D$ is strictly log-concave, as well as $f$ being log-concave and nonincreasing. By Theorem 3 in Zou (2025) (p. 5), a set of sufficient conditions for the composite function $H=D\circ\phi_{\Gamma(\hat{c})}$ to be strictly log-concave is (i) $D$ is strictly log-concave and decreasing; (ii) $\phi_{\Gamma(\hat{c})}$ is monotone and convex. We only need to show that $f$ being log-concave and decreasing leads to convex $\phi_{\Gamma(\hat{c})}$, which is true by directly verifying the second-order derivative: